# AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning

### v3

@@ Abstract
The increasing complexity of software systems and the sophistication of cyber-attacks have underscored the critical need for effective automated vulnerability detection and repair systems. Data-driven approaches using deep learning models show promise but critically depend on the availability of large, accurately labeled datasets. Yet existing datasets either suffer from noisy labels, limited range of vulnerabilities, or fail to reflect vulnerabilities as they occur in real-world software. This also limits large-scale benchmarking of such solutions. Automated vulnerability injection provides a way to directly address these dataset limitations, but existing techniques remain limited in coverage, contextual fidelity, or injection success rates. In this paper, we present AVIATOR, the first AI-agentic vulnerability injection workflow. It automatically injects realistic, category-specific vulnerabilities for high-fidelity, diverse, large-scale vulnerability dataset generation. Unlike prior monolithic approaches, AVIATOR orchestrates specialized AI agents, function agents and traditional code analysis tools, that replicate expert reasoning. It combines semantic analysis, injection synthesis enhanced with LoRA-based fine-tuning and Retrieval-Augmented Generation, as well as post-injection validation via static analysis and LLM-based discriminators. This modular decomposition allows specialized agents to focus on distinct tasks, improving robustness of injection and reducing error propagation across the workflow. Evaluations across three distinct benchmarks demonstrate that AVIATOR achieves 91%–95% injection success rates, significantly surpassing existing automated dataset generation techniques in both accuracy and scope of software vulnerabilities.
@@ I Introduction
The rapid growth in software complexity, coupled with the sophistication of cyber-attacks, poses a significant threat to the global security and stability of digital infrastructures. In 2024 alone, the total number of publicly reported vulnerabilities rose by 25% [cyberThreatIndex_2024]. Software vulnerabilities refer to weaknesses in system security requirements, design, implementation, or operation, that could be accidentally triggered or intentionally exploited, resulting in a violation of the system’s security policy [sate6_delaitre2023]. These trends highlight the urgent need for scalable and effective automated vulnerability detection and repair systems. Traditional methods, such as static program analysis, have played an important role. However, their reliance on manually crafted heuristics and the combinatorial complexity of their analysis makes them fundamentally limited, with persistent challenges in scalability, adaptability, and false alarm rates [Not_using_Static_tools, CastleBenchmarking_dubniczky2025].
@@
In response, AI-based approaches for code vulnerability detection [DLvulnDetectStudy_chakraborty2020, llmVulnSurvey_zhou2024, VulDeePecker_Li2018, SySeVR_Li2022, Rahman_2024, DiverseVul_yizheng2023, PrimeVul_ding2024] and repair [deepcodeaifixfixing_berabi2024, vulnfixvisiontransformer_fu2025, codesecurityvulnerabilityrepair_islam2024] have recently attracted considerable attention in the research community. These data-driven techniques, show strong potential in identifying and addressing a wide range of vulnerabilities. However, their effectiveness heavily relies on the quality and scale of the training data [PrimeVul_ding2024, vgx_nong2024]. This dependency has become even more critical with the emergence of Large Language Models (LLMs) with billions of parameters to train. The scarcity of high-quality, large-scale datasets therefore remains a central obstacle to advancing automated software security [PrimeVul_ding2024, llmVulnSurvey_zhou2024, DLvulnDetectStudy_chakraborty2020]. Existing vulnerability datasets can be grouped into the following categories based on how the code samples were collected and annotated:
  • Synthetic datasets: Both the code and its vulnerability labels are generated artificially, typically using predefined insecure coding patterns. A notable example is the Juliet test suite [juliet2017] which is a collection of small synthetic test cases, each demonstrating a specific weakness. It is part of a larger SARD dataset [nist_sard_2018] which contains test cases ranging from small synthetic code to production software.
    These datasets offer clean and controllable examples but often lack the structural and semantic complexity of real-world code. As a result, AI models trained on such data overfit to superficial patterns—such as specific variable names or function structures—rather than learning the underlying causes of vulnerabilities [DLvulnDetectStudy_chakraborty2020].
  • Large realistic datasets: Several widely used large-scale vulnerability datasets, including ReVeal [DLvulnDetectStudy_chakraborty2020], BigVul [BigVul_jiahao2020], CVEfixes [CVEfixes_guru2021], CrossVul [CrossVul_georgios2021] and Diversevul [DiverseVul_yizheng2023] rely on heuristic-based labeling functions. These approaches typically involve identifying vulnerability-fixing commits from public security databases, such as the National Vulnerability Database (NVD) [nist_nvd_2022], and labeling the pre-commit versions of affected functions as vulnerable. While these datasets contain code samples from real-world production software, the reliance on heuristics introduces noise and label inaccuracies. Studies have reported label accuracies as low as 25% for some datasets [PrimeVul_ding2024] [DataQuality_croft2023]. Additionally, datasets like D2A [d2adatasetbuilt_zheng2021] and Draper [draper_russell2018] use static analyzers to label code, which results in numerous false positives due to the limitations of static analysis tools. The presence of noisy or incorrect labels in all of these datasets poses challenges for training reliable vulnerability detection models [DiverseVul_yizheng2023] [PrimeVul_ding2024].
  • Reliable labels datasets: Some datasets attempt to improve label accuracy through manual expert annotation or by enforcing very strict heuristics. For instance, SVEN [SVEN_he2023] manually labels 1,606 functions, which ensures clean labels, but covering only 9 Common Weakness Enumeration (CWE) types. PrimeVul [PrimeVul_ding2024] applies rigorous heuristics, claiming human-level labeling accuracy across  7,000 vulnerable and  229,000 benign functions spanning over 140 CWEs. Yet, manual curation is inherently tedious and produces only small datasets, while overly strict heuristics hinder scalability.
  • Vulnerability injection-based datasets: Several approaches for injecting bugs into production software have been proposed and implemented [pewny2016evilcoder, dolan2016lava, roy2018bug, kashyap2019automated]. Methods such as VULGEN [vulgen_nong2023], VinJ [vinj_nong2024] and VGX [vgx_nong2024] combine pattern mining with deep learning-based localization to identify injection sites and apply vulnerability-inducing edits. These approaches scale better and produce realistic samples, but they remain constrained in both coverage and reliability. In practice these, injected bugs are limited to single-statement edits and may lack the semantic and contextual richness of real-world vulnerabilities. Moreover, the accuracy of injected datasets remains a concern: for example, VULGEN achieves only a 69% success rate in generating truly vulnerable samples, indicating potential noise and limited reliability.
@@
To address these challenges, we introduce a novel AI agentic workflow designed to automatically inject realistic, diverse, category-specific vulnerabilities for large-scale generation of high-quality vulnerability datasets. This framework, called AVIATOR (AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning) coordinates multiple AI agents - simulating expert reasoning - alongside functional agents and traditional code analysis tools. It was designed by understanding the full process of manual injection and splitting it into elementary subtasks, relying on tools and self-correction loops to provide optimal guidance to LLMs. Its design was validated by three cybersecurity experts (including two co-authors). To further boost performance, it leverages Retrieval-Augmented Generation (RAG) [rag_lewis2021] for contextual grounding and employs Low-Rank Adaptation (LoRA) [lora_hu2021] for efficient model fine-tuning. Crucially, AVIATOR requires no pattern mining, freeing it from the rigid constraints of mined edits and allowing scalable generation of vulnerabilities that span more CWE categories, with greater semantic and contextual realism.
@@
To evaluate the performance of AVIATOR in introducing vulnerabilities we conducted an experimental study on code samples from three different benchmarks used in the literature [nist_sard_2018, formai_tihanyi2023, PrimeVul_ding2024]. We leveraged formal verification tool ESBMC [esbmc_gadelha2018] for supported vulnerabilities in smaller functions and manual verification by experienced computer security researchers for the rest. We compare AVIATOR with existing automated large-scale realistic software vulnerability dataset creation technique. The results suggest a potential for the AVIATOR to outperform other techniques with regard to accuracy of dataset labels. It achieved between 91% and 95% success rates in injecting vulnerabilities across the different benchmarks.
AVIATOR is a generic framework capable of supporting injection across a large CWE spectrum. Our current instantiation already covers 133 CWE types, far exceeding VGX (23 CWEs) and demonstrating the diversity potential of datasets generated with AVIATOR. Unlike prior injection-based approaches, AVIATOR does not rely on language-specific parsers or handcrafted pattern mining. Its agentic design instead leverages LLM-based reasoning making it both broadly applicable and adaptable across programming languages.
Our contributions can be summarized as follows:
@@
  • Agentic Injection Workflow: To the best of our knowledge, we are the first to explore vulnerability injection using an AI-agentic workflow. We carefully designed a multi-agent pipeline that replicates the reasoning processes of cybersecurity experts to generate plausible vulnerability injections for large-scale quality dataset.
  • Contextual Augmentation: By integrating Retrieval-Augmented Generation modules, we ensure that vulnerability transformations are grounded in diverse, real-world code contexts.
  • Efficient Fine-Tuning: We utilize Low-Rank Adaptation (LoRA) to fine-tune LLMs on small-scale datasets with minimal computational cost, thereby enhancing scalability.
  • Empirical Validation: Our preliminary evaluation demonstrates that AVIATOR achieves between 89% and 95% success in generating plausible vulnerabilities, facilitating the creation of large-scale, accurately labeled, and diverse datasets Through our experiments, we have established the defining characteristics of high-quality software vulnerability datasets, essential for both training AI models and benchmarking purposes.
@@ II Background and Motivation
In this section, we introduce the fundamental concepts and challenges associated with software vulnerability injection to motivate our proposed approach. We first describe the properties necessary for constructing effective datasets to train and benchmark AI-based vulnerability detection and repair systems. Then, we formalize and illustrate the process of vulnerability injection in source code. Additionally, we clarify the scope and objectives that underpin our proposed solution.
@@ II-A Properties of High-Quality Vulnerability Datasets
Effective training and rigorous benchmarking of AI-based vulnerability detection and repair methods critically depend on the quality of the underlying datasets. To clearly define the desired dataset characteristics, we introduce the ABCD criteria, similar to [sate6_delaitre2023] and [dolan2016lava]:
  • Accurate: Labels must reflect ground truth with clear information on the type of vulnerability, specifically CWE identifiers and, when possible, precise vulnerability locations—essential for reliable model training and benchmark evaluation.
  • Big: The dataset should be large enough to support training of deep models and enable statistically meaningful benchmarking of detection and repair systems.
@@
  • Each sample should contain a single vulnerability type to avoid ambiguity during training and evaluation.
  • Paired benign and vulnerable samples, are ideal for repair tasks and contrastive learning.
@@ II-B Motivation
Creating datasets that rigorously fulfill the ABCD criteria and structural requirements remains an open challenge due to significant limitations inherent to existing approaches [llmVulnSurvey_zhou2024]. Synthetic datasets, while accurate and easily scalable, lack credibility and offer limited diversity, as they rely on simplistic, predefined vulnerability patterns [DLvulnDetectStudy_chakraborty2020]. Heuristic-based labeling, in contrast, provides credible and diverse samples by mining real-world open-source repositories, but it suffers from a trade-off between accuracy and scalability: stricter heuristics yield smaller but cleaner datasets [PrimeVul_ding2024], while relaxed heuristics introduce noise [DiverseVul_yizheng2023]. Static analysis-based labeling can generate large datasets, yet accuracy remains a concern due to high false-positive rates. Manual labeling, though highly accurate, is not scalable for large datasets [automaticdatalabelingsoftware_le2024].
Vulnerability injection presents a promising alternative. By design, this data augmentation technique enables scalability through automation and preserves credibility when applied to real-world codebases. However, the effectiveness of this approach depends heavily on the injection technique’s ability to introduce diverse and accurate vulnerabilities. Current state-of-the-art methods remain limited in this regard: recent studies report coverage of only 19 distinct CWE classes and achieve an injection success rate of just 69% [vulgen_nong2023, vinj_nong2024]. Such limitations constrain both the diversity and reliability of resulting datasets, particularly for benchmarking purposes. Nonetheless, these initial results suggest that with more refined techniques, vulnerability injection has the potential to systematically produce accurate, realistic, diverse, and large-scale datasets that better support the development and benchmarking of modern AI-based vulnerability detection and repair systems, as well as traditional static analyzers.
@@ II-C Vulnerability Injection
    $\tau_{i}(c^{b}_{(i)})=\langle t^{(i)}_{1},\dots,t^{(i)}_{m^{(i}}\rangle,\quad m^{(i)}\geqslant 1.$
@@ III Proposed Framework
Existing vulnerability injection techniques continue to face fundamental limitations in both coverage and reliability. These methods generally support only a limited range of vulnerability types and exhibit low success rates in injection. This is largely due to their inability to adapt to the structural and contextual intricacies of real-world code found in production, as well as a lack of deep semantic reasoning. To overcome these limitations, the proposed AVIATOR framework in figure 1, organizes the injection process as an AI-agentic workflow, explicitly mirroring the reasoning steps that a human security analyst would take.
@@
Figure 1: Architecture of the AVIATOR framework illustrating the two core modules. The top module represents the Vulnerability Injection component, which utilizes RAG and supervised fine-tuning with LoRA for guiding code transformations. The bottom module depicts the Validation component, employing LLM-based discriminators and static analysis to verify the presence and accuracy of injected vulnerabilities.

### v4

@@ Abstract
The increasing complexity of software systems and the sophistication of cyber-attacks have underscored the critical need for reliable automated software vulnerability detection. Data-driven approaches using deep learning models show promise but critically depend on the availability of large, accurately labeled datasets. Yet existing datasets either suffer from noisy labels, limited vulnerability coverage, or fail to reflect vulnerabilities as they occur in real-world software. This also limits large-scale benchmarking of such solutions. Automated vulnerability injection provides a way to address these limitations, but existing techniques remain limited in coverage, contextual fidelity, or injection success. In this paper, we present AVIATOR, the first AI-agentic vulnerability injection framework. AVIATOR decomposes vulnerability injection into a coordinated workflow of specialized AI agents, tool-based analysis, and iterative self-correction, explicitly mirroring expert reasoning. It integrates retrieval-augmented generation and lightweight LoRA-based fine-tuning to produce realistic, category-specific vulnerabilities without relying on handcrafted patterns. Across three benchmarks, AVIATOR achieves high injection fidelity (91–95%) surpassing existing injection techniques in both accuracy and vulnerability coverage. When used for data augmentation to train deep learning-based vulnerability detection (DLVD) models, AVIATOR provides the strongest downstream gains in vulnerability detection. Across models and base datasets, AVIATOR improves average F1 scores by +22% over no augmentation, +25% over VGX, holding the prior best injection success rate, and +3% over VulScribeR, the prior state-of-the-art LLM-based injection model, with +7% higher recall and no precision loss. Its augmented data exhibits the lowest distributional distortion and scales efficiently with <<2% syntax rejection at 4.3× lower cost than VulScribeR.
@@ I Introduction
The rapid growth in software complexity, coupled with the sophistication of cyber-attacks, poses a significant threat to the global security and stability of digital infrastructures. In 2024 alone, the total number of publicly reported vulnerabilities rose by 25% [8]. Software vulnerabilities refer to weaknesses in system security requirements, design, implementation, or operation, that could be accidentally triggered or intentionally exploited, resulting in a violation of the system’s security policy [11]. These trends highlight the urgent need for scalable and effective automated vulnerability detection and repair systems. Traditional methods, such as static program analysis, have played an important role. However, their reliance on manually crafted heuristics and the combinatorial complexity of their analysis make them fundamentally limited, with persistent challenges in scalability, adaptability, and false alarm rates [26, 14].
@@
In response, AI-based approaches for code vulnerability detection [4, 55, 31, 30, 42, 7, 12] and repair [1, 18, 25] have recently attracted considerable attention in the research community. These data-driven techniques show strong potential in identifying and addressing a wide range of vulnerabilities. However, their effectiveness heavily relies on the quality and scale of the training data [12, 36]. This dependency has become even more critical with the emergence of Large Language Models (LLMs) with billions of parameters to train. For vulnerability detection in particular, the practical objective is not only to curate accurate labels, but to generate data that measurably improves downstream deep learning-based vulnerability detection (DLVD) under realistic test distributions. Yet high-quality vulnerable samples remain scarce [12, 55, 4], and existing datasets often trade off realism, label fidelity, and coverage, limiting downstream generalization.
Existing vulnerability datasets can be grouped into the following categories based on how the code samples were collected and annotated:
  • Synthetic datasets: Suites such as Juliet [39] provide clean, controllable examples, but they rarely reflect the structural and semantic complexity of production code. Models trained on such data often learn superficial cues rather than vulnerability causes [4].
  • Large realistic datasets: Datasets such as ReVeal [4], BigVul [17], CVEfixes [3], CrossVul [35], and DiverseVul [7] rely on heuristic labeling from vulnerability-fixing commits and security databases (e.g., NVD [33]). While they capture real code, heuristic labeling introduces substantial noise, with reported label accuracies as low as 25% for some datasets [12, 9]. Static-analyzer-labeled datasets (e.g., D2A [54], Draper [46]) suffer similarly from false positives due to tool limitations.
  • Reliable-label datasets: Manual annotation (e.g., SVEN [21]) improves label fidelity but remains expensive and small-scale. Conversely, strict heuristics (e.g., PrimeVul [12]) improve reliability to human-level labeling accuracy, but can constrain scalability and coverage. Neither route offers a robust path to producing large, high-quality vulnerable data at scale.
  • Vulnerability injection-based datasets: Injection aims to create vulnerable samples by introducing realistic weaknesses into existing code bases, which scales better than manual curation and typically yields more natural code than purely synthetic generators [41, 13, 44, 27, 10]. Recent systems such as VULGEN [37], VinJ [38], and VGX [36] combine pattern mining with deep learning-based localization to select injection sites and apply vulnerability-inducing edits. While these approaches can generate large volumes of samples, they remain constrained in both coverage and reliability. In practice, their edits often reduce to single-statement transformations and a limited set of Common Weakness Enumeration (CWE) types, which restricts semantic and contextual realism and can lead to brittle downstream benefits. VGX [36] reports higher injection success (about 90%) but supports only 23 CWE types and relies on a partly manual pattern-mining pipeline. In parallel, VulScribeR [10] leverages LLMs with RAG and explicitly optimizes for downstream DLVD performance rather than maximizing injection fidelity; it reports an estimated injection success of 80%, and its generation quality can still depend on model variability and post-filtering. Overall, existing injection methods either achieve high success under narrow, template-like edits (e.g., VGX) or improve downstream metrics without directly targeting high-fidelity, broad-coverage injection (e.g., VulScribeR).
@@
To address these challenges, we introduce a novel AI agentic workflow designed to automatically inject realistic, diverse, category-specific vulnerabilities for large-scale generation of high-quality vulnerability datasets. This framework, called AVIATOR (AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning) coordinates multiple AI agents, simulating expert reasoning, alongside functional agents and traditional code analysis tools. It was designed by studying the full process of manual injection and splitting it into elementary subtasks, relying on tools and self-correction loops to provide optimal guidance to LLMs. Its design was validated by three cybersecurity experts (including two co-authors). To further boost performance, it leverages Retrieval-Augmented Generation (RAG) [29] for contextual grounding and employs Low-Rank Adaptation (LoRA) [23] for efficient model fine-tuning. Crucially, AVIATOR requires no pattern mining, freeing it from the rigid constraints of mined edits and allowing scalable generation of vulnerabilities that span more CWE categories, with greater semantic and contextual realism.
@@
To evaluate AVIATOR’s ability to introduce vulnerabilities, we conducted an experimental study on code samples from three widely used benchmarks [34, 48, 12]. We used the formal verification tool ESBMC [20] to validate supported vulnerabilities in smaller, self-contained functions, and we relied on manual evaluation by experienced security researchers for larger functions with missing project-level context, where ESBMC is not applicable. We compare AVIATOR with existing automated large-scale realistic software vulnerability dataset creation techniques. The results show that AVIATOR achieves highest injection fidelity, with success rates between 91% and 95% across the evaluated benchmarks.
Beyond injection fidelity, AVIATOR also delivers the strongest downstream gains under the same DLVD augmentation protocol as VulScribeR [10]. With 5k injected samples, AVIATOR improves average F1 score across models and datasets by +21.8% over no augmentation and by +24.6% over VGX, and it outperforms the prior injection-based downstream state of the art VulScribeR by +2.9% in average F1 score while increasing recall by +7.4%. Notably, these recall gains do not come at the expense of precision, which is consistent with cleaner injected vulnerabilities and a more learnable training signal. We further attribute this robustness to distributional fidelity: among all augmentation strategies, AVIATOR induces the smallest shift from the original real-world dataset (lowest KL divergence on CWE count and entropy distributions), indicating that it enriches training data without skewing it toward narrow templates. Finally, unlike VulScribeR, which explicitly optimizes downstream detection performance rather than injection correctness, AVIATOR targets correct vulnerability instantiation and validation, making it suitable not only for augmentation but also for constructing benchmark datasets with reliable labels.
[…]
The key contributions of this work are:
@@
  • AVIATOR: Agentic Vulnerability Injection Framework. To the best of our knowledge, this is the first work to explore vulnerability injection using an AI-agentic workflow. We carefully designed a multi-agent pipeline, leveraging tool-augmented subtasks with self-correction and validation loops, that replicates the reasoning processes of cybersecurity experts.
  • High-Fidelity, Broad-Coverage Vulnerability Injection. Our evaluation demonstrates that AVIATOR achieves between 91% and 95% success in generating plausible vulnerabilities across three diverse benchmarks. AVIATOR is designed as a generic framework capable of supporting injection across a wide spectrum of weakness categories; our current instantiation supports 133 CWE types, far exceeding state-of-the-art systems like VGX. This demonstrates a strong potential for constructing diverse, high-coverage datasets.
  • Downstream Effectiveness. Following the VulScribeR augmentation protocol, adding 5,000 AVIATOR-generated samples yields the strongest average downstream performance among all evaluated strategies. AVIATOR improves the average F1-score by +21.8% over no augmentation and +24.6% over VGX. Notably, it surpasses the LLM-based state-of-the-art VulScribeR by +2.9% in average F1 score and +7.4% in recall, achieving these gains without degrading precision.
[…]
@@ II Background and Motivation
In this section, we introduce the fundamental concepts and challenges associated with software vulnerability injection to motivate our proposed approach. We first describe the properties necessary for constructing effective datasets to train and benchmark Deep Learning-based vulnerability detection systems. Then, we formalize and illustrate the process of vulnerability injection in source code. Additionally, we clarify the scope and objectives that underpin our proposed solution.
@@ II-A Properties of High-Quality Vulnerability Datasets
Effective training and rigorous benchmarking of DLVD methods critically depend on the quality of the underlying datasets. To clearly define the desired dataset characteristics, we introduce the ABCD criteria, similar to [11] and [13]:
  • Accurate: Labels must reflect ground truth with clear information on the vulnerability type (CWE identifiers) and, when possible, precise vulnerability locations, to support more reliable training and faithful evaluation.
  • Big: The dataset should be large enough to support training of deep learning models and enable statistically meaningful benchmarking of detection and repair systems.
@@
  • Each sample should ideally contain a single vulnerability type to reduce ambiguity during training and evaluation.
  • Paired benign and vulnerable samples, are valuable for repair tasks and contrastive learning.
@@ II-B Motivation
Creating datasets that rigorously fulfill the ABCD criteria and structural requirements remains an open challenge due to significant limitations inherent to existing approaches [55]. Synthetic datasets, while accurate and easily scalable, lack credibility and offer limited diversity, as they rely on simplistic, predefined vulnerability patterns [4]. Heuristic-based labeling, in contrast, provides credible and diverse samples by mining real-world open-source repositories, but it exposes a persistent trade-off between accuracy and scalability: stricter heuristics yield smaller but cleaner datasets [12], while relaxed heuristics introduce noise [7]. Static analysis-based labeling can generate large datasets, yet accuracy remains a concern due to high false-positive rates. Manual annotation, though highly accurate, is not scalable for large datasets [28].
Vulnerability injection presents a promising alternative. By design, this data augmentation technique enables scalability through automation and preserves credibility when applied to real-world codebases. In principle, injection can also provide paired benign/vulnerable samples with explicit CWE labels and locations. However, the effectiveness of this approach depends heavily on the ability of the injection technique to introduce diverse and accurate vulnerabilities.
Existing injection systems have clear limitations. Pattern mining approaches such as VulGen and VinJ combine mined edit templates with deep learning-based localization to select injection sites in real code, which improves realism and scalability, but they still operate under constrained edit patterns and limited coverage, and produce noisy labels [37, 38]. VGX reports higher injection success, yet it remains tied to a partly manual pattern mining pipeline and focuses on single-statement edits with limited CWE support (23 CWE types), which narrows the space of vulnerabilities it can realistically generate [36]. In parallel, VulScribeR leverages LLMs and RAG to optimize downstream DLVD performance, but it does not explicitly target injection correctness and therefore provides noisier training signal and weaker guarantees for benchmark construction where label fidelity matters [10]. Overall, prior work tends to optimize either injection success under narrow templates (e.g., VGX) or downstream metrics without prioritizing injection fidelity (e.g., VulScribeR), leaving open the problem of achieving high fidelity injection, broad coverage, and strong downstream utility at scale.
@@ II-C Vulnerability Injection
    $\tau_{i}(c^{b}_{(i)})=\langle t^{(i)}_{1},\dots,t^{(i)}_{m^{(i)}}\rangle,\quad m^{(i)}\geqslant 1.$
@@ III Proposed Framework
Existing vulnerability injection techniques trade off fidelity, coverage, and downstream utility. Pattern-mining methods such as VulGen, VinJ, and VGX achieve scalable injection but are constrained to narrow edit templates and limited CWE coverage, while recent LLM-based approaches such as VulScribeR prioritize downstream DLVD performance without explicitly enforcing injection correctness. To address these limitations, we propose AVIATOR (Figure 1), which formulates vulnerability injection as an AI-agentic workflow that explicitly mirrors the structured reasoning process of a human security analyst.
@@
Figure 1: Architecture of the AVIATOR framework illustrating the two core modules. The top module represents the Vulnerability Injection component, which utilizes RAG and supervised fine-tuning with LoRA for guiding code transformations. The bottom module depicts the Validation component, combining LLM-based discriminators with static analysis to verify vulnerability instantiation and trigger bounded self-correction. The full framework contains 13 agents including 9 AI agents and 4 tool agents.